james drives 5 miles per hour slower than ken. If James travels 100 miles in the same time that ken travels 110 miles, find each rate. solution?
Answer by checkley77(12844)

D=RT
100=(R-5)T
OR T=100/(R-5) FOR JAMES
110=RT
OR T=110/R FOR KEN
SET THESE 2 EQUATIONS EQUAL & SOLVE FOR R.
100/(R-5)=110/R CROSS MULTIPLY,
110(R-5)=100R
110R-550=100R
100R-100R=550
10R=550
R=550/10
R=55 MPH FOR KEN.
55-5=50 MPH IS THE RATE FOR JAMES.
110=55*T
T=110/55
T=2 HOURS.
PROOF:
100=50*2
100=100